OpenStudy (anonymous):

What is the inverse of the function f(x) = x-9

OpenStudy (anonymous):

Try writing it as y = x - 9 Then swap the variables: x = y - 9 And solve for y. This will generate your inverse function.

OpenStudy (ksaimouli):

swith the values y as x and x as y and solve for y